Find the area bounded by y = 2; y = √x + 2; and y = 2 - x

1
Expert's answer
2021-05-10T18:22:41-0400

Let

"y = 2 \qquad \cdots (i)\\
y = \sqrt{x} + 2 \qquad \cdots (ii)\\
y = 2 - x \qquad \cdots (iii)"


We proceed to check the boundary points of the curve.

From (i) and (ii)


"2=\sqrt x+2\\
\sqrt x = 2-2\\
\sqrt x =0\\
x=0"

From (i) and (iii)

"2=2-x\\
x=2-2\\
x=0"

From (ii) and (iii)


"\sqrt x + 2 = 2-x\\
\implies x=0"

Since the value of x is zero for all the three functions, then, the curves not bounded and therefore, NO SOLUTION


A graph of this is shown below: